OpenStudy (anonymous):

Need help with asymptote. I am trying to solve for my Vertical Asymptote but i need help factoring: x^2-2x-2

OpenStudy (cwrw238):

this does not factor

OpenStudy (anonymous):

oh, i kinda figured that :P. Then how am i supposed to get my vertical asymptote?

OpenStudy (cwrw238):

you can find the values of x which make x^2 - 2x - 2 = zero by using the quadratic root formula or by completing the square

OpenStudy (anonymous):

ok after i set it to zero what do i have to do?

OpenStudy (ranga):

x^2-2x-2 = 0 complete the square: (x-1)^2 - 1 - 2 = 0 (x-1)^2 = 3 take square root on both sides and solve for x.

OpenStudy (anonymous):

Ok i figured that out its x=4 and x=-2, but how am i going to use that in the asymptote?

OpenStudy (anonymous):

the question was x-2/x^2-2x-2